OpenStudy (anonymous):

Mike is six less than three times his daughter's age. Mike is 36; how old is his daughter?

OpenStudy (anonymous):

Let his daughter's age be x. Mike's age = 3x - 6 3x - 6 = 36 3x = 36 + 6 = 42 x = 42/3 = 14 Mike's age = 3x - 6 = 42 - 6 = 36

OpenStudy (jiteshmeghwal9):

Let mike's daughter's age. then three times of her age=3x A/q, 3x-6=36 3x=36+6 3x=42 x=42/3=14

OpenStudy (jiteshmeghwal9):

hence, daughter's age=>x =14